Montclair Company is considering a project that will require a $660,000 loan. It presently has total liabilities of $140,000, and total assets of $700,000.
  
1. Compute Montclair’s (a) present debt-to-equity ratio and (b) the debt-to-equity ratio assuming it borrows $660,000 to fund the project.
Answer #1

Ans 1 a) 0.25

Equity = Total Assets - Total Liabilities

                           = 700000 - 140000

                          = 560000

debt-to-equity ratio = Total Liabilities / Equity

                              = 140000 / 560000

                             = 0.25

b) 1.43

debt-to-equity ratio = Total Liabilities / Equity

                              = (140000 + 660000) / 560000

                             = 1.43